# Fractured Glass, Failing Cameras: Simulating Physics-Based Adversarial Samples for Autonomous Driving Systems

### v2

@@ Abstract
While much research has recently focused on generating physics-based adversarial samples, a critical yet often overlooked category originates from physical failures within on-board cameras—components essential to the perception systems of autonomous vehicles. Camera failures, whether due to external stresses causing hardware breakdown or internal component faults, can directly jeopardize the safety and reliability of autonomous driving systems. Firstly, we motivate the study using two separate real-world experiments to showcase that indeed glass failures would cause the detection based neural network models to fail. Secondly, we develop a simulation-based study using the physical process of the glass breakage to create perturbed scenarios, representing a realistic class of physics-based adversarial samples. Using a finite element model (FEM)-based approach, we generate surface cracks on the camera image by applying a stress field defined by particles within a triangular mesh. Lastly, we use physically-based rendering (PBR) techniques to provide realistic visualizations of these physically plausible fractures. To analyze the safety implications, we superimpose these simulated broken glass effects as image filters on widely used open-source datasets: KITTI and BDD100K. We then evaluate the detection failure rates for critical classes using two most prominent object detection neural networks (CNN-based - YOLOv8 and Faster R-CNN) and Pyramid Vision Transformers. To further investigate the distributional impact of these visual distortions, we compute the Kullback-Leibler (K-L) divergence between three distinct data distributions, applying various broken glass filters to a custom dataset (captured through a cracked windshield), as well as the KITTI and Kaggle cats and dogs datasets. The K-L divergence analysis suggests that these broken glass filters do not introduce significant distributional shifts. Our goal is to provide a robust, physics-based methodology for generating adversarial samples that reflect real-world camera failures, with the overarching aim of improving the resilience and safety of autonomous driving systems against such physical threats.
@@ Introduction
Combining our rendering approach with two popular open source datasets - KITTI (Geiger et al. 2013) and BDD100k (Lin et al. 2014), we are able to generate adversarial images efficiently. A common process for testing the generated adversarial images is to find the number of false positives/negatives across the image space. However, in our case, due to the adversarial effect being local, we cannot rely simply on an image based measure. We therefore, use the adversarial images (similar to the lower left figure of Fig. 2) and extract the objects which lie within the region where the fracture exists using the ground truth bounding boxes. We then utilize YOLOv8 and Faster R-CNN (Ren et al. 2016) to find the percentage of objects that fail when the adversarial filters are applied. We also provide ablation studies to understand the distributional differences between the three set of images: Real broken glass images collected experimentally, real broken glass images collected online and the generated images. We compute the Kullbeck-Liebler (K-L) divergence for these image distributions to prove similarity of the generated images to the real broken glass images. We utilize cat images from Kaggle Cats and Dogs dataset as control to understand the difference between image distributions (PK ).
@@
where $\Omega$ is the unit hemisphere centered around surface normal n over $\omega_{i}$ such that $\omega_{i}\cdot n>0$. Abstracting the reflectance equation, we aim to create a visual render of our broken glass mesh. We have $L_{e}=0$ as glass does not emit light. Now for calculating $L_{r}$, we consider any crack between the nodes as a microfacet. Then, we can define $L_{r}$ for every crack as:
@@
While we are able to simulate realistic fractures, the primary use case for our work is to be able to generate simulated examples overlayed on existing datasets (KITTI, BDD100k) and compare them with the real on-road dataset that we created.
@@
We benchmark two types of broken glass pattern - real and simulated - on two popular open-source datasets - KITTI (Geiger et al. 2013) and BDD100k (Lin et al. 2014). The real broken glass pattern images are collected from FreePik website [^1] https://www.freepik.com and represent the baseline in our case. We collected $65$ images in total and expanded them to a set of $10,000$ images via image augmentation using random shifts, image flips and cropping techniques. We also generate $10,000$ images using our physics simulator. We then overlay these cracked glass patterns using our PBR pipeline onto every image in the KITTI and BDD100k datasets and collect the aggregate results. We use a YOLOv8 model and Faster R-CNN with pretrained weights on KITTI and BDD100k to generate object detection results.
@@
Table 1: Average precision (in percentage) of different classes in KITTI and BDD100k under different adversarial images. x provides the overlay relation between dataset and glass-crack type. Clean x Dataset - refers to directly KITTI images without any adversarial sample. RO x Dataset - refers to Real images of cracked glass collected online overlayed on clean images. Sim x Dataset - refers to simulated crack patterns overlayed on clean images.
@@
KITTI (YOLOv8) 0.25 Pedestrian 32.67 81.39 22.07
Truck 12.39 3.85 3.84
Car 69.6 61.5 67.2
0.5 Pedestrian 25.64 69.72 17.84
@@
Table 1 shows the results of the average precision (AP) under the adversarial images generated using the two types of crack patterns (collected online and simulated) for different classes. For KITTI, the AP of other classes drop as expected with the decrease in AP corresponding to the percentage of image occupied with the truck class recording the highest drop. For BDD100k, we aggregated the AP for the super - categories: person, vehicles and food. This is because a lot of objects in BDD100k occupy smaller area in the image frame making it difficult to get meaningful results from all categories.
A very intriguing result is that the pedestrian class has a multifold increase in AP under the real broken glass patterns. A similar pattern is also observed in the vehicle class (BDD100k with Faster R-CNN). While this trend might seem counter-intuitive, it resonates with the results in Fig. 2 where the confidence of the car increases because of an edge. This in fact shows that the AP is highly dependent on the crack pattern making it extremely important to create defense methodologies to mitigate these adversarial attacks.
@@
Figure 6: K-L divergence of different pairs of image distributions. The three different datasets are: RC - Real on-road dataset (see Fig. 2), KITTI and Cats. The two different broken glass filters are: RO - Real (collected online) and Sim - Simulated. The different bars present the K-L divergence between: A - (Sim x KITTI) vs (Sim x KITTI); B - (Sim vs RO); C - (Clean RC vs KITTI); D - (Broken RC) vs (RO x KITTI); E - (Broken RC) vs (Sim x KITTI); F - KITTI vs Cats (x provides the overlay relation between dataset and filter type).
@@ Appendix B Static Experiment
Appendix C Dynamic Experiment
@@
Appendix D Real glass fracture images
@@
Figure 15: Top - Inference of YOLOv8 of glass fracture overlaid on KITTI. Bottom - Inference of clean image for comparison.

### v3

@@ Abstract
While much research has recently focused on generating physics-based adversarial samples, a critical yet often overlooked category originates from physical failures within on-board cameras—components essential to the perception systems of autonomous vehicles. Camera failures, whether due to external stresses causing hardware breakdown or internal component faults, can directly jeopardize the safety and reliability of autonomous driving systems. Firstly, we motivate the study using two separate real-world experiments to showcase that indeed glass failures would cause the detection based neural network models to fail. Secondly, we develop a simulation-based study using the physical process of the glass breakage to create perturbed scenarios, representing a realistic class of physics-based adversarial samples. Using a finite element model (FEM)-based approach, we generate surface cracks on the camera image by applying a stress field defined by particles within a triangular mesh. Lastly, we use physically-based rendering (PBR) techniques to provide realistic visualizations of these physically plausible fractures. To assess the safety implications, we apply the simulated broken glass effects as image filters to two autonomous driving datasets- KITTI and BDD100K- as well as the large-scale image detection dataset MS-COCO. We then evaluate detection failure rates for critical object classes using CNN-based object detection models (YOLOv8 and Faster R-CNN) and a transformer-based architecture with Pyramid Vision Transformers. To further investigate the distributional impact of these visual distortions, we compute the Kullback-Leibler (K-L) divergence between three distinct data distributions, applying various broken glass filters to a custom dataset (captured through a cracked windshield), as well as the KITTI and Kaggle cats and dogs datasets. The K-L divergence analysis suggests that these broken glass filters do not introduce significant distributional shifts. Our goal is to provide a robust, physics-based methodology for generating adversarial samples that reflect real-world camera failures, with the overarching aim of improving the resilience and safety of autonomous driving systems against such physical threats.
Codehttps://github.com/manavprabhakar/camera-failure
@@ Introduction
Combining our rendering approach with three popular open source datasets - KITTI (Geiger et al. 2013), BDD100k (Yu et al. 2020) and MS-COCO (Lin et al. 2014), we are able to generate adversarial images efficiently. A common process for testing the generated adversarial images is to find the number of false positives/negatives across the image space. However, in our case, due to the adversarial effect being local, we cannot rely simply on an image based measure. We therefore, use the adversarial images (similar to the lower left figure of Fig. 2) and extract the objects which lie within the region where the fracture exists using the ground truth bounding boxes. We then utilize YOLOv8, Faster R-CNN (Ren et al. 2016) and Pyramid Vision Transformer (PVTv2) (Wang et al. 2022) to find the percentage of objects that fail when the adversarial filters are applied. We also provide ablation studies to understand the distributional differences between the three set of images: Real broken glass images collected experimentally, real broken glass images collected online and the generated images. We compute the Kullbeck-Liebler (K-L) divergence for these image distributions to prove similarity of the generated images to the real broken glass images. We utilize cat images from Kaggle Cats and Dogs dataset as control to understand the difference between image distributions (PK ).
@@
where $\Omega$ is the unit hemisphere centered around surface normal n over $\omega_{i}$ such that $\omega_{i}\cdot n>0$.
Abstracting the reflectance equation, we aim to create a visual render of our broken glass mesh. We have $L_{e}=0$ as glass does not emit light. Now for calculating $L_{r}$, we consider any crack between the nodes as a microfacet. Then, we can define $L_{r}$ for every crack as:
@@
While we are able to simulate realistic fractures, the primary use case for our work is to be able to generate simulated examples overlayed on existing datasets (KITTI, BDD100k, MS-COCO) and compare them with the real on-road dataset that we created.
@@
We benchmark two types of broken glass pattern - real and simulated - on three popular open-source datasets - KITTI (Geiger et al. 2013), BDD100k (Yu et al. 2020) and MS-COCO (Lin et al. 2014). The first two represent specific autonomous driving domain while the last one is a general purpose image dataset. The real broken glass pattern images are collected from FreePik website [^1] https://www.freepik.com and represent the baseline in our case. We collected $65$ images in total and expanded them to a set of $10,000$ images via image augmentation using random shifts, image flips and cropping techniques. We also generate $10,000$ images using our physics simulator. We then overlay these cracked glass patterns using our PBR pipeline onto every validation image in the datasets and collect the aggregate results. We use three model architectures YOLOv8, Faster R-CNN and PVTv2 model with pretrained weights to generate object detection results.
@@
Table 1: Average precision (in percentage) of different classes in KITTI, BDD100k and MS-COCO under different adversarial images. x provides the overlay relation between dataset and glass-crack type. Clean x Dataset - refers to directly the particular images without any adversarial sample. RO x Dataset - refers to Real images of cracked glass collected online overlayed on clean images. Sim x Dataset - refers to simulated crack patterns overlayed on clean images.
@@
KITTI (YOLOv8) 0.5 Pedestrian 25.64 69.72 17.84
@@
MS-COCO (Faster R-CNN) 0.5 Person 0.035 0.024 0.024
Vehicles 2.14 1.45 1.87
Food 35.34 28.07 30.65
0.75 Person 0.032 0.022 0.023
Vehicles 1.56 1.05 1.07
Food 24.59 18.85 22.00
@@
Table 1 shows the results of the average precision (AP) under the adversarial images generated using the two types of crack patterns (collected online and simulated) for different classes. For KITTI, the AP of other classes drop as expected with the decrease in AP corresponding to the percentage of image occupied with the truck class recording the highest drop. For BDD100K with PVTv2-B0, we see that the drop in AP is largest in the simulated images but overall, the trend is maintained with the pedestrian class showing the steepest drop. For MS-COCO, we aggregated the AP for the super - categories: person, vehicles and food. This is because a lot of objects in MS-COCO occupy smaller area in the image frame making it difficult to get meaningful results from all categories.
Figure 6: K-L divergence of different pairs of image distributions. Datasets: RC - Real on-road dataset (see Fig. 2), KITTI and Cats. Filters: RO - Real (collected online) and Sim - Simulated. K-L divergence between (x - overlay relation): A - (Sim x KITTI) vs (Sim x KITTI); B - (Sim vs RO); C - (Clean RC vs KITTI); D - (Broken RC) vs (RO x KITTI); E - (Broken RC) vs (Sim x KITTI); F - KITTI vs Cats.
A very intriguing result is that the pedestrian class has a multifold increase in AP under the real broken glass patterns. While this trend might seem counter-intuitive, it resonates with the results in Fig. 2 where the confidence of the car increases because of an edge. This in fact shows that the AP is highly dependent on the crack pattern making it extremely important to create defense methodologies to mitigate these adversarial attacks.
@@
In this work, we address black-box adversarial attacks stemming from real-world, naturally occurring physical phenomena, not artificially crafted to exploit specific model vulnerabilities. We assume no knowledge of the model attributes, weights or architecture, ensuring attacks are transferable across various models. Physical adversarial methods (Translucent Patch, RP2) can all be termed as occlusions of either the camera or the objects being captured. The adversariality comes from the effect of the model inference due to these occlusions. Our PBR pipeline blends the cracks with source images as translucent, blurry patterns, impacting latent space encoding rather than causing direct occlusion, resulting in incorrect detections.
@@ Appendix B Static Experiment
Figure 13: (a) Top - detections on a clean image; bottom - detections on an adversarial image.(b) YoLo fails to detect the person (c) Glass cracks allows the model to detect the person.
Appendix C Increased AP for pedestrians in KITTI
We would like to point out that the increased AP for the pedestrian class was something that even we were surprised at first. However, a careful-qualitative deep-dive analysis helped us understand that this was occurring as a result of the glass cracks making it easier for the model to classify pedestrians because of enhanced edges around them. This wasn’t an edge artifact but instead the glass crack acting as an additional edge boundary clearly separating the pedestrian and the background. A similar result was also observed in [1] where the overall AP was increased in adversarial images.
Appendix D Dynamic Experiment
@@
Appendix E Real glass fracture images
@@
Figure 16: Top - Inference of PTv2 on a clean image from BDD100k. Bottom - Inference for a real broken glass image overlaid on BDD100k for comparison. We see two extra false positives in on the right side (truck, car) and several false negatives for the pedestrian class on the left of the adversarial image.